OpenStudy (anonymous):

Please help write out this problem!!!!!!!!!! The total value of a collection of quarters and nickels is $3.75. If the total number of coins is 27, how many of each type of coins are there?

OpenStudy (anonymous):

(.05)(N)+(.25)(Q)27=3.75?

OpenStudy (foolaroundmath):

Why did you multiply Q with 27? Reason that out for me.

OpenStudy (anonymous):

I'm not sure where to put the 27

OpenStudy (foolaroundmath):

If you have N nickels, what is the value of N nickels?

OpenStudy (anonymous):

.05

OpenStudy (foolaroundmath):

One nickel is 0.05, what is the value of N of them?

OpenStudy (anonymous):

5N?

OpenStudy (foolaroundmath):

0.05*N

OpenStudy (foolaroundmath):

Similarly one quarter is 0.25, the value of Q quarters is ____.

OpenStudy (foolaroundmath):

Yes, now what is the total value of N nickels and Q quarters? (in terms of N and Q) What should this be equal to (as given in the question)

OpenStudy (foolaroundmath):

Yes, so 0.05*N + 0.25*Q = 3.75 Now it is given that the number of nickels + the number of quarters is 27.

OpenStudy (foolaroundmath):

So, can you make another equation relating Q and N with 27?

OpenStudy (anonymous):

.05N+.25Q-27=3.75??

OpenStudy (foolaroundmath):

Number of nickels + number of quarters = 27 N + Q = 27 and the other equation was 0.05*N+0.25*Q=3.75 Can you solve these two?

OpenStudy (anonymous):

.25Q+.05(27-N)=3.75?????????????????????????????????????????????????????????????????

OpenStudy (foolaroundmath):

0.25Q+0.05(27-Q)=3.75 (Since N+Q = 27, N = 27-Q and then substitute in the second equation)